OpenStudy (mathhelp346):

PLEASE HELP: Trevor is a salesperson who earns a weekly salary and a commission that is 7% of his weekly sales. In one week Trevor's sales were $952.00 and his weekly income was $466.64. In another week his sales were $2515.00 and his weekly income was $576.05. a. Write a linear equation in slope-intercept form for Trevor's weekly income, y, in terms of his weekly sales, x. b. What is Trevor's weekly salary?

OpenStudy (anonymous):

Still there?

OpenStudy (mathhelp346):

yea

OpenStudy (anonymous):

ok, we can do this (me AND you).

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

This is just a problem in getting a line from 2 points. y = mx + b. Y will be income. b will be his weekly salary. x will be sales. weekly salary will a constant.

OpenStudy (mathhelp346):

ok will there be two equations?

OpenStudy (anonymous):

You are given 2 values for y and 2 values for x. You can then use (y2 - y1)/(x2 - x1) = m. You can get m from the % of sales commission. Only one equation constructed from the 2 points.

OpenStudy (mathhelp346):

so do i get multiply the weekly sales by .07 first?

OpenStudy (mathhelp346):

then i do y2-y1/x2-x1

OpenStudy (anonymous):

You multiply mx by x to get the commission in $.

OpenStudy (mathhelp346):

but i need the coordinates to do y2-y1/x2-x1

OpenStudy (anonymous):

You might want to start with figuring out his base salary. That part will be very easy because you only need one point for that since so much info is given in the problem statement. Use y - mx = b. Use either point, and you will get the rsame correct answer. I worked it out both ways and the answer is the same.

OpenStudy (mathhelp346):

y-mx=b?

OpenStudy (anonymous):

You were given the coordinates. If you read my posts and ascribe the corresponding x's and y's, you will have your coordinates. As i said we are BOTH doing the problem. I do not hand out answers. I give strong hints and set up the problem for people.

OpenStudy (mathhelp346):

so so the coordinates would be (952,466.64) (2515, 576.05)? but do you multiply the weekly sales(952,2515) by .07 first to get his commission?

OpenStudy (mathhelp346):

would x be the weekly sales or his commission

OpenStudy (anonymous):

You're doing well. You have now identified the coordinates and that is very good. You would multiply m by x (commission rate by sales) to get the commission. They already gave you the y (total income for the week), so you just have to subtract mx from y to get b or the base salary. You can use either point, they both give the same answer.

OpenStudy (anonymous):

I have already identified x for you. If you still have a question on that, go back and read my posts.

OpenStudy (mathhelp346):

ok so i find the slope and then substitute it into y=mx+b to get the salary?

OpenStudy (anonymous):

I'm going to answer your questions by asking you questions. What is y?

OpenStudy (mathhelp346):

total income for the week

OpenStudy (anonymous):

good. What makes up the total income for the week (in words and variables)?

OpenStudy (mathhelp346):

b(weekly sales)

OpenStudy (mathhelp346):

x(sales)

OpenStudy (mathhelp346):

b is weekly salary

OpenStudy (anonymous):

You're close. You're right about x being sales, but you're off on b. b is his salary. Salary + sales commission (sales x commission rate or x multiplied by m) = income = y. With that info, you are 90% done.

OpenStudy (mathhelp346):

so if i solve 466.64=.07(952)+b i will get the salary?

OpenStudy (mathhelp346):

but wouldnt i have 2 equations since i can sub in either one? which one do i use?

OpenStudy (anonymous):

yes. If you used the other point, you would still get the same salary also. You can use either point.

OpenStudy (mathhelp346):

so do i write down both equations?

OpenStudy (anonymous):

They are just 2 points on the same line. You don't even need the second point to do this problem. It's there for 2 reasons. 1) to give a person practice digging out of a word problem only that which is essential (sufficient and necessary) to do the problem. And 2) so you can confirm your answer (equation and value for salary) with the second point, which again, isn't that important. You would need the second point if you were not given the value for m, sales commission.

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

Last time. One equation. I cannot help you any further. Now, 98% of the work is done.

OpenStudy (mathhelp346):

ok thanks

OpenStudy (anonymous):

good luck.